### v2

@@ 5 Conclusion
This paper is devoted to establishing a series of sufficient and equivalent conditions of the robust ICKKTM for spectral norm regularized convex matrix optimization problems. We extend the results for nuclear norm regularized convex optimization problems in [16] to spectral norm regularized convex optimization problems.There is still much work to be done, such as how to apply the variational properties of proximal mapping of spectral norm function to directly prove the equivalence between (a) and (c) in Theorem 4.4, whether the conclusions in this paper or in [16] can be extended to the Ky Fan $k$-norm regularized problems, or whether the variational properties of the spectral norm function obtained in this paper can be applied to provide stability results for the generalized variational inequality problem with the spectral norm function. These are all the topics we will study in the future.

### v3

@@ 5 Conclusion
This paper is devoted to establishing a series of equivalent conditions of the robust ICKKTM for spectral norm regularized convex matrix optimization problems. We extend the results for nuclear norm regularized convex optimization problems in [16] to spectral norm regularized convex optimization problems. There is still much work to be done, such as how to apply the variational properties of proximal mapping of spectral norm function to directly prove the equivalence between (a) and (c) in Theorem 4.4, whether the conclusions in this paper or in [16] can be extended to the Ky Fan $k$-norm regularized problems, or whether the variational properties of the spectral norm function obtained in this paper can be applied to provide stability results for the generalized variational inequality problem with the spectral norm function. These are all the topics we will study in the future.